AUTOIONIZATION OF HIGH ORBITAL ANGULAR MOMENTUM STATES OF CALCIUM MONOFLUO-RIDE

\underline{JEFFREY J. KAY}, DANIEL S. BYUN, JONATHAN R. BARCHI, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

The use of a double-resonance laser excitation scheme and proper choice of intermediate electronic state has enabled the observation of high orbital angular momentum states of calcium monofluoride with ℓ = 3 and one or two quanta of core vibrational excitation. These states undergo vibrational autoionization by Δv^{+} = 1 and Δv^{+} = 1 processes. The structure of these states and their observed rates of autoionization are compared with the predictions of the long-range force model of core-nonpenetrating states.